# Diffusion Forcing for Multi-Agent Interaction Sequence Modeling

### v1

@@ Abstract
Understanding and generating multi-person interactions is a fundamental challenge with broad implications for robotics and social computing. While humans naturally coordinate in groups, modeling such interactions remains difficult due to long temporal horizons, strong inter-agent dependencies, and variable group sizes. Existing motion generation methods are largely task-specific and do not generalize to flexible multi-agent generation. We introduce MAGNet ( Multi-Agent Diffusion Forcing Transformer ), a unified autoregressive diffusion framework for multi-agent motion generation that supports a wide range of interaction tasks through flexible conditioning and sampling. MAGNet performs dyadic prediction, partner inpainting, and full multi-agent motion generation within a single model, and can autoregressively generate ultra-long sequences spanning hundreds of v. Building on Diffusion Forcing, we introduce key modifications that explicitly model inter-agent coupling during autoregressive denoising, enabling coherent coordination across agents. As a result, MAGNet captures both tightly synchronized activities (e.g., dancing, boxing) and loosely structured social interactions. Our approach performs on par with specialized methods on dyadic benchmarks while naturally extending to polyadic scenarios involving three or more interacting people, enabled by a scalable architecture that is agnostic to the number of agents. We refer readers to the supplemental video, where the temporal dynamics and spatial coordination of generated interactions are best appreciated. Project Page
@@ 1 Introduction
Understanding and generating multi-person interactions is a fundamental challenge in computer vision and graphics with applications in robotics, virtual reality, and social computing. The requirements for such generative models vary significantly: robots need to react to human motion, artists may want flexible control over interacting motion through keyframing, and virtual agents may need to populate virtual worlds by generating long natural social interactions from minimal conditioning.
Moreover, social situations often involve more than dyadic (two-agent) interactions. Despite its importance, existing methods are typically designed for specific dyadic tasks—such as reaction synthesis given the other agent’s motion or joint prediction of both agents—requiring different models for different scenarios.
In this work, we introduce Multi-Agent Diffusion Forcing Transformer (MAGNet), a unified framework for modeling and generating multi-agent interactions. MAGNet generates realistic sequences for both synchronized interactions such as dancing and boxing as well as asynchronous social interactions, and can be autoregressively rolled out over time. Thanks to its transformer architecture, it naturally accommodates varying numbers of agents, and enables flexible sampling strategies at inference time to support multiple tasks including: Partner prediction and Partner in-painting–predicting the motion of an agent given partial or complete motion of other agents, Joint Future Prediction–predicting future motion of all agents given past motion context, and more. Furthermore, our model can operate in an agentic manner [cen2025readytoreact], where it runs independently on each agent to generate motion from that agent’s perspective based on observed actions of others, enabling distributed inference on individual robots or autonomous virtual agents.
Our approach is inspired by recent advances in autoregressive diffusion, particularly Diffusion Forcing (DF) [chen2025diffusion] and TEDi [shi2024tedi], which have shown promising results for sequence modeling through token-wise noise scheduling in video generation and single-person motion modeling. We propose a transformer-based diffusion model trained over sequences of tokens, where each token represents an agent at a specific timestep and receives a different noise level. This simple approach learns not only the joint distribution over all tokens, but also the conditional distributions over any subsequence—a key property identified by DF that enables flexible inference. This flexibility is particularly valuable for multi-agent modeling, enabling a single model to perform multiple tasks. For example, to generate reactive motion to another agent at test time, the model takes in clean motion tokens of the conditioning agent while denoising the missing agent’s motion tokens. Similarly, joint generation of all agent motion can be achieved by keeping past motion tokens clean and denoising future motion tokens of all agents in an autoregressive manner.
To effectively model multi-agent interactions, we introduce key design changes. Unlike single-person motion generation where motion can be modeled in isolation, multi-person interactions fundamentally depend on inter-agent relationships—how agents are positioned and oriented relative to each other. We therefore represent all motion, including each agent’s global trajectory, in relative transformations between agents grounded in per-frame canonical frames. This makes the model agnostic to absolute positioning in the world space. Second, we learn a discrete latent space for the motion of each agent. Third, we model the dynamics between agents with Diffusion Forcing Transformer (DFoT) using motion tokens that encode each agent’s latent pose information as well as the relative transformation to all other agents.
With a single unified architecture, we achieve strong performance across multiple downstream tasks including dyadic interaction generation, and partner inpainting and prediction, while naturally handling polyadic scenarios with three or more people. Notably, we demonstrate continuous dyadic social interactions without requiring text conditioning, relying instead on the motion dynamics themselves. Our method performs on par with specialized prior works in quantitative metrics, while generating diverse, long interaction sequences. While Figure 1 illustrates key results, we strongly encourage viewing the result videos , as the temporal dynamics and spatial coordination are best appreciated in motion.
@@ 2 Related Work
Single-Agent Motion Generation
@@
We focus on methods that generate the whole body motion of one or more people, for methods that predict root trajectories of pedestrians, see a recent survey [huang2025vision]. Single-agent motion generation has evolved from RNNs and VAEs  [Fragkiadaki2015RecurrentNM, julieta2017motion, aksan, maluleke2024synergy] to diffusion models [dabral2023mofusion, shi2024amdm] and transformers[tevet2022mdm]. While promising, these methods struggled with long-term temporal coherence and high-quality motion synthesis. Diffusion models marked a breakthrough in motion generation. Text-conditioned methods like MDM [tevet2022mdm] and MoFusion [dabral2023mofusion] generate diverse, temporally consistent human motion, while transformer-based approaches such as T2M-GPT [zhang2023t2mgpt] use self-attention to capture long-range dependencies. More recently, TEDi [shi2024tedi] advanced long sequence generation through temporally-entangled diffusion that recursively denoises a motion buffer, enabling arbitrary-length sequences without stitching artifacts. However, these methods remain limited to single-agent scenarios.
Dyadic Motion Generation
@@
Dyadic motion generation methods model two-person interactions often by modeling inter-agent dependencies with cross-attention or diffusion. Text-conditioned methods like InterGen generate synchronized two-person motions from text descriptions [liang2024intergen], while ExPI predicts future motions by modeling dependencies among agents’ past trajectories [guo2022multiperson]. ReMoS and ReGenNet apply diffusion models with spatio-temporal cross-attention for partner inpainting—synthesizing one agent’s motion conditioned on another’s complete motion sequence [ghosh2024remos, xu2024regennet]. However, these approaches are primarily unidirectional, meaning they lack the mechanisms to treat generated outputs as reciprocal feedback that can dynamically influence other agents. ARFlow introduces a multi-modal diffusion framework for both partner inpainting and prediction, though limited to short clips [jiang2025arflow]. Human-X employs autoregressive diffusion for low-latency generation in VR/AR [ji2025humanx], but is optimized for real-time reactive motion and cannot produce long-horizon coordinated behaviors. Music-conditioned approaches target dance generation. Duolando combines GPT architectures with off-policy reinforcement learning for music-conditioned partner inpainting [siyao2024duolando]. DuetGen and Dyadic Mamba focus on choreography-driven dance synthesis [ghosh2025duetgen, tanke2025dyadicmamba]. Methods like BUDDI [mueller2023buddi], Reaction Priors [Fang_2024_CVPR], and Ponimator [liu2025ponimator] learn dyadic human priors to reconstruct two people from images or video.
@@
Closest to our approach is Ready-to-React, which unifies vector quantization, diffusion, and autoregressive generation for partner prediction in an agentic manner—each agent can independently run a model for reactive motion generation [cen2025readytoreact]. However its architecture is restricted to agentic sampling and cannot handle joint future prediction. A fundamental limitation across these methods is their inability to scale beyond dyadic interactions. The cross-attention mechanisms in Interformer [guo2023interformer], ReMoS [ghosh2024remos], and ReGenNet [xu2024regennet] are designed to attend from one agent to one other agent, making it unclear how to extend them when multiple other agents are present. Other methods’ architectures similarly assume two-agent scenarios. Table 1 provides a systematic comparison across five key capabilities essential for comprehensive multi-agent motion generation. Our approach can naturally handle more than two agents by adding more agent motion tokens, while unifying these diverse tasks within a single framework.
@@ 3 Method
We introduce the Multi-Agent Diffusion Forcing Transformer (MAGNet), a unified autoregressive diffusion framework designed for flexible motion generation among multiple interacting agents. Our approach is inspired by the Diffusion Forcing Transformer (DFoT) [song2025historyguidedvideodiffusion]. We design the motion tokens to be multi-component, encoding the single-agent latent pose alongside the necessary pairwise inter-agent transforms to all partners. Below we discuss how we represent the inter-agent relationships, latent motion encoding, and the multi-agent diffusion forcing transformer.
@@
Figure 3: Multi-Agent Diffusion Forcing Transformer (MAGNet). Left (Training): Each agent’s motion is encoded by a VQ-VAE into latent pose tokens, forming motion tokens $m_{i}^{p}$ by appending latent vectors with transform parameters. Tokens from all agents are interleaved and processed by a Diffusion Forcing Transformer with independently noised tokens. Right (Inference): The model enables flexible conditioning: known (blank) tokens are fixed, while unknown tokens are causally denoised. This supports partner in-painting, joint prediction, and agentic turn-taking, where agents alternately generate motion and highlighted streams can run independently (e.g., on separate robots).
@@ 3.1 Motion Representation
Our goal is to model a sequence of $T$ time steps involving $P$ interacting people. Each person’s motion is represented via their body shape, $\mathbf{\beta}\in\mathbb{R}^{P\times 10}$, and their joint rotations, $\Theta_{t}=[\theta^{0}_{t},...,\theta^{J}_{t}],\Theta_{t}\in\mathbb{R}^{T\times P\times J\times 6}$, which can be mapped to a 3D mesh with $J$ joints using a human body model [Pavlakos2019_smplifyx]. We use the 6D rotation representation for joint rotations.
To better capture spatial relations, we employ two coordinate frames and the transforms between them: (1) a root frame at the pelvis and (2), following EgoAllo  [yi2025egoallo]) a per–time step canonical frame obtained by projecting the root frame onto the floor plane Figure 2. This representation makes the model invariant to absolute world positioning while avoiding dependence on a fixed initial reference frame. Notably such representations prevent translation coordinate magnitudes from growing during long generation. We denote rigid transforms by $\mathbf{T}^{X\to Y}_{t}$ from $X$ frame to $Y$ frame at time step t and parameterize each in 9D (6 for rotation, 3 for translation). In particular,
$\mathbf{T}^{\text{can}\to\text{root}}_{t}\in\mathbb{R}^{T\times P\times 9}$ (1)
defines the “canonical $\to$ root transform” at time step $t$.
Let $\Delta{t=1}$ denote one-step transitions in the canonical frame. The intra-agent temporal transform from t1t\!-\!1 to $t$ is
@@
Building on prior work, we introduce pairwise inter-agent transforms at each time step:
@@
which encodes the transform from self to each partner in the canonical frame at time $t$.
3.2 Latent Motion Encoding
@@
We adopt a latent-space approach, widely used to improve stability in long-horizon synthesis [chen2023executing, rombach2021highresolution]. We first train a conditional VQ-VAE [van2017neural] restricted to the single-agent pose and orientation information while inter-agent and temporal relations are modeled by DFoT.
VQ-VAE Encoding. We compress the input $x_{t}=(\Theta_{t},\mathbf{T}^{\text{can}\to\text{root}}_{t})$ into a latent representation conditioned on $c_{t}=(\beta,\Delta\mathbf{T}^{\text{can}}_{t})$. With temporal stride $\omega$ ($\omega=4$), the encoder yields token embeddings:
$\mathbf{H}=\mathrm{Enc}(x_{1:T}\mid c_{1:T})=[\,\mathbf{h}_{1},\ldots,\mathbf{h}_{T/\omega}\,],\quad\mathbf{h}_{i}\in\mathbb{R}^{d}.$ (4)
The quantizer maps each $h_{i}$ to the nearest codebook vector:
$\mathbf{Z}=[\mathbf{z}_{1},\mathbf{z}_{2},...,\mathbf{z}_{T/\omega}],\quad\mathbf{z}_{i}\in\mathbb{R}^{d}.$ (5)

### v2

@@ Abstract
Understanding and generating multi-person interactions is a fundamental challenge with broad implications for robotics and social computing. While humans naturally coordinate in groups, modeling such interactions remains difficult due to long temporal horizons, strong inter-agent dependencies, and variable group sizes. Existing motion generation methods are largely task-specific and do not generalize to flexible multi-agent generation. We introduce MAGNet (Multi-Agent Generative Network), a unified autoregressive diffusion framework for multi-agent motion generation that supports a wide range of interaction tasks through flexible conditioning and sampling. MAGNet performs dyadic and polyadic prediction, partner inpainting, partner prediction, and agentic generation all within a single model, and can autoregressively generate ultra-long sequences spanning hundreds of motion steps. We explicitly model inter-agent coupling during autoregressive denoising, enabling coherent coordination across agents. As a result, MAGNet captures both tightly synchronized activities (e.g., dancing, boxing) and loosely structured social interactions. Our approach performs on par with specialized methods on dyadic benchmarks while naturally extending to polyadic scenarios involving three or more interacting people. Please watch the supplemental video, where the temporal dynamics and spatial coordination of generated interactions are best appreciated. Project Page.
@@ 1 Introduction
Understanding and generating multi-person interactions is a fundamental challenge in computer vision and graphics, with applications spanning robotics, virtual reality, and social computing. Requirements are diverse: robots must react to human motion in an online streaming fashion, artists may want flexible control through keyframing, and virtual agents must produce long, natural social interactions from minimal conditioning. Moreover, social situations often extend beyond dyadic pairs to polyadic groups of three or more.
This diversity of requirements has led to a fragmented landscape of approaches. Most methods are tailored to specific dyadic tasks— either reacting to an agent motion, or joint prediction of both agents—requiring separate models for each scenario. Architecturally, prior work either encodes inter-agent relationships through cross-attention between fixed agent pairs [9, 26, 12], limiting generalization to arbitrary group sizes, or restricts inference to agentic-only sampling modes [2], sacrificing flexibility. No single existing model supports the full spectrum of multi-agent generation tasks.
We introduce Multi-Agent Generative Network (MAGNet), a unified framework that addresses this gap. MAGNet is a transformer-based diffusion model designed around a relational motion tokenization that flexibly supports a wide range of motion generation tasks while naturally scaling beyond two agents.
Rather than representing motion in a global frame or as a fixed-size concatenation over agents, we construct agent–time motion tokens: each token corresponds to one agent at one timestep and is built to expose both what the agent does and how it relates to others. Concretely, a token contains a compact motion representation (via a VQ-VAE latent pose) together with explicit pairwise relative transforms to all other agents, grounded in per-frame local canonical frames. On top of this representation, we train with diffusion forcing [3, 25], assigning each token an independent noise level. This enables a single model to learn conditional distributions over arbitrary subsets of agents and timesteps, supporting partner prediction, inpainting, joint future prediction, and rollout within the same architecture (e.g., reactive synthesis by keeping partner tokens clean while denoising the target agent; joint rollout by keeping past tokens clean while denoising the future across agents).
Our work makes three key contributions. First, we design a relational, world-agnostic multi-agent motion representation: by embedding pairwise transforms in per-frame canonical coordinates, interactions are parameterized by who relates to whom rather than absolute positions, making the model agnostic to the number of participants and enabling polyadic scenarios ($P>2$) without architectural modification. Specifically, we eliminate the reliance on cross-attention to inject inter-agent information. Second, unlike prior diffusion forcing formulation where only a single token is denoised at a time, in multi-agent generation,
agents at the same timestep must remain coordinated even when their tokens are corrupted with different noise levels. Pairing each agent’s VQ-VAE latent with explicit pairwise transforms keeps inter-agent geometry available under any noise configuration, allowing tokens to be denoised independently, conditioned on observed agents. This design also enables agentic inference [2] for distributed deployment. Third, without any text conditioning, MAGNet is competitive with specialized baselines on partner prediction and inpainting, and substantially improves the harder regimes of dyadic generation and long-horizon synthesis—e.g., improving FD by 89% over Ready-to-React [2]—while scaling seamlessly to polyadic interactions. Our code will be released. Please see Fig. 1 and the supplementary video.
@@ 2 Related Work
Single-Agent Motion Generation.
@@
We focus on methods that generate the whole body motion of one or more people, for methods that predict root trajectories of pedestrians, see a recent survey [13]. Single-agent motion generation has evolved from RNNs and VAEs  [8, 19, 1, 18] to diffusion models [5, 24] and transformers[29]. While promising, these methods struggled with long-term temporal coherence and high-quality motion synthesis. Diffusion models marked a breakthrough in motion generation. Text-conditioned methods like MDM [29] and MoFusion [5] generate diverse, temporally consistent human motion, while transformer-based approaches such as T2M-GPT [35] use self-attention to capture long-range dependencies. More recently, TEDi [25] advanced long sequence generation through temporally-entangled diffusion that recursively denoises a motion buffer, enabling arbitrary-length sequences without stitching artifacts. However, these methods remain limited to single-agent scenarios.
Dyadic Motion Generation.
@@
Dyadic motion generation methods model two-person interactions often by modeling inter-agent dependencies with cross-attention or diffusion. Text-conditioned methods like InterGen generate synchronized two-person motions from text descriptions [16], while ExPI predicts future motions by modeling dependencies among agents’ past trajectories [11]. ReMoS and ReGenNet apply diffusion models with spatio-temporal cross-attention for partner inpainting—synthesizing one agent’s motion conditioned on another’s complete motion sequence [9, 32]. However, these approaches are primarily unidirectional, meaning they lack the mechanisms to treat generated outputs as reciprocal feedback that can dynamically influence other agents. ARFlow introduces a multi-modal diffusion framework for both partner inpainting and prediction, though limited to short clips [15]. Human-X employs autoregressive diffusion for low-latency generation in VR/AR [14], but is optimized for real-time reactive motion and cannot produce long-horizon coordinated behaviors. Music-conditioned approaches target dance generation. Duolando combines GPT architectures with off-policy reinforcement learning for music-conditioned partner inpainting [26]. DuetGen and Dyadic Mamba focus on choreography-driven dance synthesis [10, 28]. Methods like BUDDI [21], Reaction Priors [7], and Ponimator [17] learn dyadic human priors to reconstruct two people from images or video.
@@
Closest to our approach is Ready-to-React, which unifies vector quantization, diffusion, and autoregressive generation for partner prediction in an agentic manner—each agent can independently run a model for reactive motion generation [2]. However its architecture is restricted to agentic sampling and cannot handle joint future prediction. A fundamental limitation across these methods is their inability to scale beyond dyadic interactions. The cross-attention mechanisms in Interformer [12], ReMoS [9], and ReGenNet [32] are designed to attend from one agent to one other agent, making it unclear how to extend them when multiple other agents are present. Other methods’ architectures similarly assume two-agent scenarios. Table 1 provides a systematic comparison across five key capabilities essential for comprehensive multi-agent motion generation. Our approach can naturally handle more than two agents by adding more agent motion tokens, while unifying these diverse tasks within a single framework.
@@ 3 Method
We introduce the Multi-Agent Generative Network (MAGNet), a unified autoregressive diffusion framework for flexible motion generation among multiple interacting agents. In our formulation, each token represents a single agent at a single timestep, and the diffusion model is trained with independent per-token noise levels via diffusion forcing [3]. Each token is a composite representation encoding the agent’s latent pose alongside explicit pairwise transforms to every other agent in the scene. Below we discuss how we represent the inter-agent relationships, latent motion encoding, and the Multi-Agent Generative Network.
@@
Figure 3: Multi-Agent Generative Network (MAGNet). Left (Training): Each agent’s pose at a given time is encoded by a VQ-VAE into latent pose tokens, forming motion tokens $m_{i}^{p}$ by appending latent vectors with transform parameters. Tokens from all agents across all timesteps are denoised by a transformer with independently noised tokens. Right (Inference): The model enables flexible conditioning: known (solid) tokens are fixed, while unknown tokens are causally denoised. This supports partner inpainting, joint prediction, and agentic turn-taking, where agents alternately generate motion and highlighted streams can run independently (e.g., on separate robots).
@@ 3.1 Motion Representation
Our goal is to model a sequence of $T$ timesteps involving $P$ interacting people. Each person’s motion is represented via their body shape, $\mathbf{\beta}\in\mathbb{R}^{P\times 10}$, and their joint rotations, $\Theta_{t}=[\theta^{0}_{t},\dots,\theta^{J}_{t}],\Theta_{t}\in\mathbb{R}^{T\times P\times J\times 6}$, which can be mapped to a 3D mesh with $J$ joints using a human body model [22]. We use the 6D rotation representation for joint rotations. To capture spatial relationships while remaining invariant to absolute world positioning, we define two coordinate frames per agent: (1) a root frame at the pelvis, and (2) a per-timestep canonical frame obtained by projecting the root onto the ground plane, following [34], as shown in Figure 2. Working in canonical coordinates prevents translation magnitudes from growing unboundedly during long-horizon generation.
We denote rigid transforms by $\mathbf{T}^{X\to Y}_{t}$ from $X$ frame to $Y$ frame at timestep t and parameterize each in 9D (6 for rotation, 3 for translation). In particular, the “canonical $\to$ root transform” at timestep $t$ is defined as:
$\mathbf{T}^{\text{can}\to\text{root}}_{t}\in\mathbb{R}^{T\times P\times 9}.$ (1)
Let $\Delta\mathbf{T}$ denote one-timestep transitions in a frame. The intra-agent temporal transform in the canonical frame from t1t\!-\!1 to $t$ is
@@
We then model pairwise inter-agent transforms at each timestep:
@@
which encodes the transform from self to each partner in the canonical frame at time $t$. For a scene with $P$ agents, each agent carries P1P-1 such transforms—one to every other agent—which are concatenated into the motion token representation described in Section 3.1. This design allows the model to handle arbitrary group sizes without architectural modification
@@
We adopt a latent-space approach, widely used to improve stability in long-horizon synthesis [4, 23]. We first train a conditional VQ-VAE [30] restricted to the single-agent pose
information while inter-agent and temporal relations are modeled by DFoT.
VQ-VAE Encoding.
We compress the input $x_{t}=(\Theta_{t},\mathbf{T}^{\text{can}\to\text{root}}_{t})$ into a latent representation conditioned on $c_{t}=(\beta,\Delta\mathbf{T}^{\text{can}}_{t})$. Here, the pose $x_{t}$ comprises the joint rotations $\Theta_{t}$ and the root’s height and tilt (roll and pitch), denoted by $\mathbf{T}^{\text{can}\to\text{root}}_{t}$. With temporal stride $\omega=4$, the encoder yields token embeddings $\mathbf{H}=\mathrm{Enc}(x_{1:T}\mid c_{1:T})=[\,\mathbf{h}_{1},\ldots,\mathbf{h}_{T/\omega}\,]$, where $\mathbf{h}_{i}\in\mathbb{R}^{d}$. The quantizer then maps each $\mathbf{h}_{i}$ to the nearest codebook vector, producing $\mathbf{Z}=[\mathbf{z}_{1},\ldots,\mathbf{z}_{T/\omega}]$, where $\mathbf{z}_{i}\in\mathbb{R}^{d}$. Conditioning the VQ-VAE decoder is necessary because a person’s body shape ($\beta$) directly influences its vertical root placement, and the temporal transform ($\Delta\mathbf{T}^{\text{can}}_{t}$) provides the necessary context of direction and velocity for accurate pose reconstruction.